4. a. The exponential average formula used to predictthe length of the next CPU burst is given by the recursion equation: at, +(1-a)t, n+1 Where t denotes the nth cPU bust time, and 0sasl. Explain the roll of a in the equation. u. Supposethe five actual uru bursts ot a process in milliseconds are: to-6, t 4, t 6, t 8, t 6. And suppose the initial predicted time is to 7. Compute the next predicted time T c. What are the implications of assigning the following values to the parameters used by the algorithm? i, a 0 and T 100milliseconds ii. and 10milliseconds i
Answer #1

a)

In this formula tn is most recent information. Tn stores past history.The parameter α controles the relative weights of recent and past history in out prediction.

b)

let α=0.5

T1 = T0 * α + ( 1 - α ) * t0

T1 = (0.5 x 7) + ( (1 - 0.5) x 6 )
     = 3.5 +3.0
   = 6.5

T1 = T1 * α + ( 1 - α ) * t1
T2 = (0.5 x 6.5) + ( (1 - 0.5) x 4 )
     = 3.25 +2.0
   = 5.25

T3 = T2 * α + ( 1 - α ) * t2
T3 = (0.5 x 5.25) + ( (1 - 0.5) x 6 )
    = 2.625 +3.0
    = 5.625

T4 = T3 * α + ( 1 - α ) * t3
T4 = (0.5 x 5.625) + ( (1 - 0.5) x 8 )
    = 2.8125 +4.0
    = 6.8125

T5 = T4* α + ( 1 - α ) * t4
T5 = (0.5 x 6.8125) + ( (1 - 0.5) x 6 )
    = 3.40625 +3.0
    = 6.40625

c)

1)

if α=0

Tn+1 = α tn + (α-1) Tn

         =0*tn + (0-1)Tn

        = 0 - Tn

       = - Tn (Time always is in postive) so

    = Tn


Tn+1 = Tn =

so T0 = T1 = T2 = T3 = T4...............=Tn = 100 milliseconds

2)

if α=1

Tn+1 = α tn + (α-1) Tn

         =1 *tn + (1 -1)Tn

        = tn - 0 * Tn

       = tn

   


Tn+1 = tn

so T0 = T1 = T2 = T3 = T4...............=Tn = tn = 10 milli seconds